At 23:22 we started recording data under the new conditions.
Merger, Tape server (WS3): 1408 kB/s writing ok
Leakage current (WS4): 3.27 uA
DAQ (WS2): temperature checked, clock status:ok, adc calibration:ok, white rabbit check:failed
Online monitoring (WS 5): done |